(4a-3) = a + 13/a + 1
The solution is

Answer by Maths68(1474)

(4a-3) = a + 13/a + 1
Multiply by (a+1) both sides
(4a-3)(a+1) = (a+1)(a + 13)/(a + 1)
(4a-3)(a+1) =a + 13
4a^2+4a-3a-3=a+13
4a^2+a-3=a+13
4a^2+a-3=a+13
Subsract (a) and add (3) both sides
4a^2+a-3-a+3=a-a+13+3
4a^2=16
Divide by 4 both sides
4a^2/4=16/4
a^2=4
a^2=2^2
a=2
Check:
(4a-3) = a + 13/a + 1
Put a=2
[4(2)-3]=(2+13)/(2+1)
8-3= 15/3
5=5